# Conceptualization, Methodology, Writing – original draft, Software, Writing – review & editing

### v3

@@ 5.1.2 Segment-level
The HBSGRP model for segment-level V–V near misses introduces random parameters within a hierarchical structure (Table 7). Specifically, the intercept and three time-varying vehicle dynamics—relative speed, relative deceleration, and relative distance—were allowed to vary across segments (Fig. 5.1.2). This specification recognizes that longitudinal interactions in uninterrupted midblock contexts are shaped by both driver behavior and roadway characteristics (lane configuration, median type, driveway density). All remaining covariates—relative acceleration, jerk, heading and steering differences, traffic volume, lane attributes, and access density—were treated as fixed effects given their comparatively stable influence across segments. The HBSGRP model yields substantially better fit than the fixed-effect alternative, confirming the value of partial pooling for capturing heterogeneity across the nine directional segments.
Refer to caption Refer to caption
(a) $\beta_{\mu,0[1,18]}$ (b) $\gamma_{\mu,1[1,18]}$
Refer to caption Refer to caption
(c) $\gamma_{\mu,2[1,18]}$ (d) $\gamma_{\mu,3[1,18]}$
Refer to caption Refer to caption
(e) $\beta_{\sigma,0[1,18]}$ (f) $\gamma_{\sigma,1[1,18]}$
Refer to caption Refer to caption
(g) $\gamma_{\sigma,2[1,18]}$ (h) $\gamma_{\sigma,3[1,18]}$
(i) $\beta_{\xi,0[1,18]}$
Fig. 11.: Posterior estimates for V-V near-misses at segments
@@
    Hierarchical Bayesian Spatial Grouped Random Parameter
@@
For the location parameter ($\mu$), the random effects reveal meaningful differences across segments that align with expected driving dynamics. Relative speed shows a consistently positive effect, confirming that faster closing speeds sharply reduce 2D-TTC and intensify near-miss severity. The magnitude of this effect, however, varies across roadway contexts Fig. 5.1.2. Segments with uninterrupted high-speed flows (e.g., S11) show the strongest impacts, while access-rich segments (e.g., S6, S7, and S13) display wider credible intervals, reflecting greater variability in interactions. Relative deceleration carries a negative sign overall, but its effect is less stable, suggesting that braking sometimes helps dissipate conflicts in smoother-flow environments, whereas in higher-speed segments the same braking often signals critical near-miss events. Relative distance exhibits a consistently negative association, indicating that shorter headways reliably elevate severity by restricting the available buffer for evasive maneuvers. The posterior variation across segments suggests that spacing behavior differs between divided and undivided facilities. Finally, in some segments with sparse block maxima samples, the intercept’s credible intervals approach zero, highlighting weaker estimation stability under limited data conditions.
Among the fixed effects, relative acceleration shows a significant negative association with $\mu$, indicating that acceleration phases are systematically linked to more severe near-misses on segments. This pattern contrasts with intersections, where acceleration often reflects flow resumption under green signals. Along segments, however, aggressive acceleration during overtaking or merging reduces spacing and elevates risk. Jerk also exhibits a negative effect, underscoring how abrupt longitudinal control contributes to critical outcomes. Lateral control indicators—heading and steering differences—display positive coefficients, suggesting that lateral adjustments provide modest increases in 2D-TTC where lane-changing flexibility is available. Driveway density is positively associated with $\mu$, a result that may seem counterintuitive but likely reflects anticipatory caution by drivers in access-rich environments. In contrast, undivided medians show a strong negative effect, confirming that exposure to opposing traffic flows produces consistently more severe interactions.
For the scale parameter ($\sigma$), the random effects for relative speed, deceleration, and distance are all negative, indicating that extreme longitudinal dynamics compress outcome variability and yield uniformly critical near-miss conditions. This pattern is evident in Fig. 5.1.2, where high-speed and short-spacing conditions cluster tightly around lower 2D-TTC values, producing narrow posteriors in segments such as S12 and S15. By contrast, segments like S10 display wider credible intervals, reflecting weaker estimation stability under sparse BM samples. Among the fixed effects, jerk and heading difference increase $\sigma$, suggesting that volatile acceleration or lateral corrections introduce greater heterogeneity in near-miss outcomes. In contrast, lane number and lane width reduce $\sigma$, implying that additional maneuvering space fosters more predictable interactions. Driveway density increases dispersion, consistent with the variability created by frequent merging and crossing maneuvers. Median type again shows a strong negative effect, reinforcing that undivided facilities compress severity outcomes into consistently critical ranges.
Finally, the shape parameter ($\xi$) remains negative and statistically significant across segment groups, confirming a bounded upper tail consistent with the Weibull domain. This finding indicates physical and behavioral constraints on extreme near-miss severity, governed by speed environment, roadway geometry, and driver reaction capacity. Overall, the segment-level results emphasize the central role of longitudinal dynamics—acceleration, braking, and headway—in determining near-miss severity, while roadway features such as medians, lane width, and access density primarily shape variability. The strong performance of the HBSGRP model underscores its value in capturing these segment-specific dynamics, which fixed-parameter structures fail to represent, particularly in accounting for the spatially contingent effects of speed and spacing in midblock contexts.
@@
The HBSGRP model for segment-level V–I near-misses offers insights into how longitudinal dynamics and roadway design jointly shape the severity and variability of extreme 2D-TTC outcomes (Table 8, Fig. 5.1.2). In this model, the intercept and relative distance are specified as group-level random effects in both the location and scale components, capturing spatial heterogeneity in how spacing influences near-miss severity across segments. All other covariates, including relative speed, acceleration, deceleration, jerk, heading difference, steering angle difference, traffic volume, lane attributes, driveway density, median type, and left-lane presence—are modeled as fixed effects to represent their more consistent contributions across segment environments.
@@
    Hierarchical Bayesian Spatial Grouped Random Parameter
@@
For the location parameter, reduced clearance to roadway boundaries exerts a strong negative effect on 2D-TTC, confirming that closer spacing to infrastructure consistently produces more critical near-miss outcomes. The random effect for distance further indicates that this influence varies across segment groups, reflecting sensitivity to local geometric and operational conditions such as median presence or driveway density. Relative speed shows a notable divergence: while the fixed-parameter model suggests higher speeds intensify severity, the grouped specification produces a small positive estimate, implying that in some contexts faster travel may coincide with greater clearance from boundaries (e.g., drivers maintaining lane center on free-flow arterials). This reversal highlights how heterogeneous speed environments fundamentally alter the role of speed in shaping V–I severity. Relative acceleration is positive and significant, suggesting that acceleration phases often represent clearance from roadside constraints (e.g., moving away from curbs or barriers) rather than heightened risk. By contrast, relative deceleration retains a negative effect across both models, indicating that braking near boundaries is typically associated with abrupt evasive actions, sharply reducing 2D-TTC and producing more critical near-miss conditions.
Several geometric and operational features also shape V–I near-miss severity at the segment level. Heading and steering differences exhibit significant positive coefficients, indicating that lateral adjustments toward or away from boundaries provide modest clearance gains and mitigate severity. Traffic volume, lane number, and lane width also emerge as positive contributors in the HBSGRP specification, suggesting that wider, higher-capacity segments allow greater separation from roadside features, producing slightly larger 2D-TTC values. The effect of driveway density, however, is particularly notable: while insignificant in the fixed-parameter model, the grouped specification yields a large negative estimate with a wide CI. This suggests that dense roadside access points intensify boundary conflicts by forcing frequent merging and turning maneuvers, though the data remain sensitive to localized characteristics. Median type is also significant, with undivided facilities consistently producing lower 2D-TTC values, reflecting heightened exposure to opposing flows and limited recovery space near boundaries. The left-lane variable shows a sign reversal, with the fixed model suggesting higher severity in interior lanes and the grouped model indicating greater severity in outer lanes. This pattern highlights heterogeneity across segment groups, where lane position relative to curbs, medians, or roadside access alters the likelihood of boundary-related near misses.
The scale parameter provides complementary insights into the variability of boundary-related near misses. The random effect for relative distance is positive, indicating that reduced clearance to roadside features not only increases severity but also broadens variability, as drivers adopt divergent strategies when negotiating minimal spacing with curbs, medians, or barriers. Relative speed also shows a positive association with scale, suggesting that higher speeds amplify heterogeneity in boundary interactions—contrasting with V–V near misses, where speed reduces dispersion—because drivers respond differently to infrastructure proximity under fast-moving conditions. Acceleration, by contrast, has a negative coefficient in the grouped model, implying that acceleration phases near boundaries often reflect smoother clearance maneuvers, yielding more predictable outcomes. Relative deceleration also reduces variability, indicating that braking close to infrastructure compresses outcomes into uniformly critical events.
Geometric attributes further shape variability. Lane width exerts a positive effect, showing that while wider lanes provide greater average clearance, they also increase variability by enabling a broader range of lateral positioning relative to boundaries. The undivided median type strongly increases scale, underscoring that the absence of a raised median not only heightens risk but also expands outcome variability due to heterogeneous driver avoidance maneuvers. The left-lane variable is also positive, indicating that boundary proximity in outer lanes produces more heterogeneous near-miss outcomes compared to interior lanes. Finally, driveway density, though statistically unstable, suggests that variability is strongly site-dependent, as frequent access points create unpredictable driver–infrastructure interactions.
The shape parameter remains negative and significant across both specifications, confirming that the upper tail of the 2D-TTC distribution is bounded. This supports the Weibull domain for extremes and indicates that even in V–I segment interactions, physical and behavioral limits constrain how severe the most critical boundary-related near-misses can become. Variation in the shape parameter across segment groups underscores the role of local geometry, access density, and operational conditions in shaping the extremal behavior of these events.
Segment-level analyses show that roadway geometry and vehicle dynamics jointly shape near-miss severity and variability. For V–V events, relative speed and distance dominate—speed amplifies severity while distance compresses variability, producing more deterministic outcomes under close-following conditions. For V–I events, the effect of speed is heterogeneous, sometimes reducing severity when clearance to boundaries is maintained, but elevating it when high speeds bring vehicles closer to curbs or medians. Access-related factors, particularly driveway density and median type, are critical for V–I outcomes, underscoring the role of access management and median design in roadside safety. Across both interaction types, the negative shape parameter indicates bounded extremes, reflecting physical and behavioral limits on crash risk escalation.
@@ 6 Model validation
The study corridor dataset provided precise geolocation for each extreme near-miss event (Eqn. 32) along with annual corridor-wide crash totals. Although annual crash data offer a useful aggregate benchmark, they lack precise geolocation of individual crashes and therefore cannot be directly matched to near-miss predictions at the event level. To address this limitation, validation was instead designed as an event-level case–control exercise, assessing whether the model assigns higher exceedance probabilities to more severe near-miss events. Posterior exceedance probabilities were computed (Eqn. 30) across severity thresholds from –0.1 to –0.9 seconds, and classification performance was evaluated using ROC–AUC, a threshold-independent metric well suited to rare-event prediction.
@@
ROC curves were generated for the four HBSGRP–UGEV variants, covering both V-V and V-I interactions at the segment and intersection levels. The results (Figure 6) reveal clear distinctions across near-miss types and roadway contexts. The V–V segment model achieved the highest predictive accuracy (AUC $\approx$ 0.89), reflecting the influence of dynamic variability in driver behavior. Relative speed, deceleration, and distance emerged as significant random parameters, capturing heterogeneity in conflict severity across segments. At the same time, fixed effects such as heading difference, steering difference, jerk, lane width, and driveway density contributed additional explanatory power, indicating that both dynamic interactions and roadway design shape V–V near-miss risks. The V–I model at the segment level achieved moderately strong accuracy (AUC $\approx$ 0.79) but displayed greater sensitivity to severity thresholds. Only relative distance emerged as a significant random parameter, while other vehicle dynamics (e.g., speed, deceleration, etc.) and roadway features (e.g., lane width, driveway density, median type, and etc.) contributed as fixed effects. This suggests that V–I risks are shaped primarily by localized conditions, with dynamic variability playing a comparatively smaller role than in V–V interactions.
At intersections, the V–V model achieved moderately strong accuracy (AUC $\approx$ 0.82) with low threshold sensitivity. Random parameters for relative speed and relative distance captured cross-location heterogeneity in approach and gap-acceptance conditions, while fixed effects— including deceleration/acceleration, steering and heading differences, jerk, and a left-turn indicator—explained lateral maneuvering and control effects. Together, these terms allow the model to represent both longitudinal and lateral conflict dynamics under signalized conditions, yielding stable performance. In contrast, the intersection-level V-I model performed weakest (AUC $\approx$ 0.75) and showing large fluctuations across thresholds, which reflects the complexity of boundary-related near-misses in intersections where turning trajectories, curb proximity, traffic signals, and pedestrian activity all interact.
Taken together, the results demonstrate that the HBSGRP–UGEV framework yields robust and consistent predictive accuracy for V–V interactions, particularly along corridor segments, where conflict dynamics are more regular and less influenced by intersectional variability. By contrast, V–I models—especially at intersections—showed lower stability and greater sensitivity to threshold specification, reflecting the complexity of boundary-related near-miss shaped by turning flows, curb geometry, and multimodal activity. These findings suggest that segment-level V–V estimates are the most operationally reliable for proactive monitoring, while intersection-level V–I models may require further refinement. Incorporating richer representations of curb design, pedestrian exposure, and high-resolution V-I proximity measures could strengthen stability and predictive performance in future applications.
@@ 7 Summary and conclusions
This study addresses the persistent challenge of estimating short-term crash occurrence risk (COR) in urban corridor-wide traffic environments, where conventional surrogate safety models often overlook detailed vehicle geometry, roadway context, and site-specific behavioral variability. Existing approaches have typically focused on vehicle–vehicle (V–V) interactions and neglected vehicle–infrastructure (V–I) conflicts, limiting their applicability at the corridor scale. To overcome these gaps, we introduced a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator that incorporates vehicle dimensions, dynamic states, and roadway boundaries, enabling the detection of extreme near-miss events from both V–V and V–I interactions.
@@
Extreme events were modeled through a nonstationary Univariate Generalized Extreme Value (UGEV) distribution embedded within a Hierarchical Bayesian Spatial Grouped Random Parameter (HBSGRP) framework. This structure allowed partial pooling across intersections and directional segments while preserving local heterogeneity, addressing data sparsity and contextual variation. High-resolution vehicle trajectories from the Argoverse-2 dataset were integrated with HD maps to extract near-miss events. Vehicle dynamics (relative speed, acceleration, deceleration, jerk, heading and steering differences, and traffic volume) and roadway features (lane width, driveway density, median type, turning movements, and lane position) were incorporated as covariates, with grouped random parameters specified for key dynamic variables. Grouped random parameters were specified for temporally varying dynamics, relative speed, relative deceleration, and relative distance, to capture their context-dependent effects. For benchmarking, a baseline hierarchical Bayesian fixed-parameter (HBSFP) model was also estimated.
Results showed that the HBSGRP model consistently outperformed its fixed-parameter counterpart, yielding lower DIC, WAIC, and LOOIC values and demonstrating greater stability. Random-parameter vehicle dynamics revealed that relative speed, deceleration, and distance dominated V–V near misses at the segment level, whereas only relative distance was significant for V–I near misses. At intersections, V–V near misses were dominated by relative speed and distance, while no dynamic variables were significant for V–I near misses. Fixed covariates such as heading difference, steering difference, lane width, driveway density, and median type also showed significant effects, demonstrating the ability of the model to capture both dynamic behaviors and roadway design influences.
Validation results further confirmed model credibility. ROC–AUC validation confirmed robust predictive accuracy, with V–V segment models performing best (AUC $\approx$ 0.89), followed by intersection V–V ($\approx$ 0.82)), while V–I models showed moderate performance ($\approx$ 0.79 for segments, $\approx$ 0.75 for intersections). These findings indicate that while V–V risks can be reliably modeled, V–I interactions, particularly at intersections, remain more variable and complex.
While the proposed framework advances corridor-wide COR modeling, several aspects warrant further development. First, reliance on AV onboard sensors may introduce measurement bias in traffic state estimation and near-miss detection. Second, the 2D-TTC formulation simplifies nonlinear vehicle dynamics, which could understate risks during extreme evasive maneuvers. Third, the framework does not yet capture near-miss severity stratification. Future research offers opportunities to enrich this framework. Extending it to multivariate HBSGRP-UGEV models with complementary surrogate measures such as post-collision velocity change ($\Delta V$) would provide deeper insights into crash severity dynamics. Incorporating vulnerable road user (VRU) near-misses, validating across diverse urban networks, calibrating for transferability, and benchmarking against GPD methods will further enhance robustness and generalizability. As AV deployment expands, increasingly rich trajectory datasets will strengthen the statistical power of HBSGRP–UGEV and open new opportunities for nuanced, multimodal safety assessments. Within this evolving data landscape, the proposed framework provides a scalable and adaptable foundation for proactive corridor-wide safety monitoring. By enabling early identification of high-risk conditions and supporting real-time traffic management, the framework aligns with the broader Vision Zero agenda, advancing the long-term goal of safer, more efficient, and equitable urban mobility.

### v4

@@ 5.1.2 Segment-level
The segment-level HBSGRP model for V–V near-misses specifies grouped random parameters to capture heterogeneity across directional segments (Table 7). Random effects were assigned to the location and scale intercepts and to three time variant covariates: relative speed, relative deceleration, and relative distance, allowing their contributions to vary by segment . All remaining covariates (relative acceleration, jerk, heading and steering differences, traffic volume, lane attributes, and driveway/access density) were treated as fixed effects to maintain a parsimonious specification when additional between-segment variation was not reliably identifiable. The HBSGRP model provides a better fit than the HBSFP alternative, supporting partial pooling for modeling segment-level variation in extreme near-miss risk.
@@
    Hierarchical Bayesian Spatial Grouped Random Parameters
@@
For the segment-level V–V model, the grouped random effects indicate that closing dynamics shift the fitted extreme distribution through paired changes in the location and scale components, and that these joint shifts vary across segment contexts. Relative speed exhibits a positive association with the location component and a negative association with the scale component, implying that as closing speed increases, the fitted distribution moves toward more critical extremes while simultaneously becoming more concentrated around those extremes. The posterior variation across segments indicates that this coupled shift is strongest on segments characterized by more uninterrupted progression, whereas access-rich segments show wider posterior intervals, consistent with more heterogeneous interaction patterns. Relative distance shows a negative association with the location component and a negative association with the scale component, indicating that tighter spacing is linked to more critical tail behavior and less dispersion in extreme outcomes. Relative deceleration contributes through a coupled location/scale pattern whose magnitude is segment-dependent, consistent with the fact that braking can represent either conflict dissipation or crisis response depending on prevailing speeds and spacing; this heterogeneity is reflected in the segment-specific posterior intervals. In segments with fewer block-maxima observations, the intercept posteriors are less precise, but the corridor-level pattern remains: closing-rate and spacing jointly govern how the extreme distribution shifts and concentrates, which is the mechanism through which COR changes.
Among the fixed effects, relative acceleration and jerk have negative associations with the location component, whereas jerk has a positive association with the scale component, suggesting that aggressive longitudinal control is associated with a shift in the external distribution and increased dispersion in outcomes. Heading difference has positive associations with both the location and scale components, whereas steering difference has a positive location association but a scale association whose interval overlaps zero, indicating that lateral misalignment tends to shift the external distribution, with less consistent evidence that it systematically changes dispersion. Exposure and geometry covariates also show coupled behavior: volume and lane width have positive location associations, while lane number and lane width have negative scale associations, suggesting that cross-sectional capacity primarily manifests through changes in dispersion rather than a consistent shift in the center of extremes. Driveway density shows positive associations at both the location and scale levels, indicating that access activity is linked to both a shift and a widening of the extreme distribution. The median type (undivided) shows negative associations in both location and scale, and left-lane position shows a positive location association and a negative scale association, indicating that cross-section and lane placement shape COR through simultaneous changes in the fitted center and the dispersion of extreme near-miss outcomes. Overall, segment-level COR is best interpreted as the result of these coupled location–scale shifts, with grouped vehicle dynamics providing the primary source of spatial heterogeneity (Table 7).
Finally, the shape parameter ($\xi$) is negative and statistically significant across segment groups, indicating a bounded upper tail for extreme near-miss behavior. Overall, longitudinal dynamics (speed, braking, and spacing) drive segment-level extremes, while roadway features (median type, lane width, and driveway density) primarily modulate variability. The improved fit of the HBSGRP model shows that allowing key dynamics to vary by segment is important for capturing midblock heterogeneity that fixed-parameter models miss.
@@
The HBSGRP model for segment-level V–I near-misses offers insights into how longitudinal dynamics and roadway design jointly shape the severity and variability of extreme 2D-TTC outcomes (Table 8, Fig. 5.1.2). In particular, the location intercept is random and significant, and the relative distance is random in both components, with a negative association in the location component and a positive association in the scale component. Implying that reduced clearance shifts extremes toward more critical outcomes while increasing dispersion, and that this sensitivity varies by segment. All other covariates, including relative speed, acceleration, deceleration, jerk, heading difference, steering angle difference, traffic volume, lane attributes, driveway density, median type, and left-lane presence, are modeled as fixed effects to represent their more consistent contributions across segment environments.
@@
    Hierarchical Bayesian Spatial Grouped Random Parameters
@@
Regarding the fixed parameters, for instance Relative speed is positive in both components, indicating that higher-speed boundary interactions are associated with a systematic shift and greater dispersion. Relative acceleration is positive in the location component and negative in the scale component, whereas relative deceleration is negative in both components, indicating that braking-dominated boundary encounters are associated with a shift toward more critical extremes and a more concentrated distribution. Heading and steering differences are positive in the location component and negative in the scale component, suggesting that lateral deviations coincide with shifted extremes but with reduced dispersion.
Roadway and exposure variables show similar coupled behavior. Lane number is positive in the location component and negative in the scale component, while lane width is positive in both components, indicating that additional lanes align with reduced dispersion, whereas wider lanes align with greater dispersion. Driveway density is negative in the location component and does not show a stable scale association, indicating access-rich segments align with more critical extremes, with dispersion effects less consistent. Undivided medians are negative in the location component and positive in the scale component, linking undivided sections to more critical and more variable extremes. Vehicle position (left lane) is negative in the location component and positive in the scale component, indicating lane context shifts extremes while increasing dispersion.
The shape parameter is negative and significant across sites (Fig. 10(e)), indicating a bounded upper tail and supporting the Weibull domain for extreme 2D-TTC. This implies that, even at the segment level, physical and behavioral constraints limit how severe boundary-related extremes can become. Differences in this parameter across segments suggest that local geometry and operating conditions influence tail behavior.
Overall, segment results show that vehicle dynamics and roadway features jointly govern both the shift and spread of extremes. For V–V interactions, relative speed, relative deceleration, and relative distance are the primary drivers, jointly shifting the distribution towards more critical outcomes and concentrating it under high closing rate, hard braking, and tight spacing conditions. For V–I interactions, relative distance is the dominant dynamic driver, with its segment-specific variability reflecting how clearance to curbs/medians governs boundary-related extremes.
@@ 6 Model validation
The trajectory-based near-miss dataset provides precise geolocation for each extracted extreme interaction (Eq. 32) because all vehicle states are map-referenced and time-stamped at high frequency. These geolocated extremes can be aggregated to corridor elements (intersections and directional segments) and compared spatially. In contrast, the available historical crash records for the study corridor are reported with imprecise geolocation; therefore, they cannot be matched to specific near-miss trajectories. Consequently, direct external validation that links predicted near-miss risk to individual crash records is not feasible with the available crash data. Validation is therefore designed to assess: (i) out-of-sample discrimination on unseen extreme blocks and (ii) spatial generalizability through site-by-site comparisons of expected versus observed extremes with uncertainty. Both HBSFP–UGEV and HBSGRP–UGEV were evaluated using these validation procedures; only the superior-performing HBSGRP–UGEV results are reported.
Model performance was evaluated using an out-of-sample event case–control design. The dataset was split at the block level into 70% training and 30% testing. For each severity threshold $\omega\in{-0.1,-0.2,\ldots,-0.9}$ sec, a test-set block was labeled as severe if its observed block maximum exceeded $\omega$. For each held-out block, the posterior exceedance probability was computed from the fitted HBSGRP–UGEV and HBSFP-UGEV model using Eq. (30). Because the HBSGRP–UGEV specification consistently outperformed the HBSFP–UGEV baseline in predictive criteria and posterior stability, only HBSGRP–UGEV validations are reported here. Receiver operating characteristic (ROC) curves were constructed by varying the classification probability cutoff, and predictive performance was summarized using the area under the ROC curve (ROC-AUC), which is threshold-independent and suitable for rare-event classification.
@@
ROC curves for the four HBSGRP–UGEV variants (V–V and V–I; segments and intersections) are shown in Fig. 6. The V–V segment model achieved the strongest identification (AUC $\approx 0.89$), followed by the V–V intersection model (AUC $\approx 0.82$). The V–I models exhibited lower identification, with AUC $\approx 0.79$ for segments and AUC $\approx 0.75$ for intersections. These results indicate that the framework more reliably distinguishes severe V–V near-misses, particularly along midblock segments, than infrastructure-related V–I near-misses in intersection environments, where turning trajectories, curb/median proximity, and signalized operations introduce greater variability.
While ROC–AUC assesses overall discrimination, it does not indicate whether the model reproduces spatial variation across corridor elements. To evaluate spatial generalizability, expected versus observed extremes were compared site by site, accounting for uncertainty. For each site $k$ and threshold $\omega$, the observed number of severe extremes in the test set was computed as the count of held-out blocks at site $k$ whose observed block maximum exceeded $\omega$. The corresponding expected number of severe extremes was computed by summing the model-based exceedance probabilities across the same held-out blocks. Uncertainty in expected counts was quantified using posterior predictive simulation, propagating posterior uncertainty in exceedance probabilities to replicated exceedance counts and summarizing the resulting predictive intervals. The expected-observed results indicate that the fitted models generally reproduce the spatial ordering of extremes across corridor elements, supporting the ability of the HBSGRP–UGEV framework to capture site-to-site variation rather than only aggregate frequency. The comparisons also reveal a systematic calibration pattern across thresholds: the model tends to overestimate exceedances at more severe thresholds (i.e., deeper tail levels) and underestimates them at less severe thresholds. This behavior suggests that the fitted tail response may be overly sensitive for the most extreme conditions and comparatively conservative for milder near-miss conditions.
Together, the hold-out ROC–AUC results (Fig. 6) and the site-by-site expected–observed comparisons with uncertainty provide complementary evidence of predictive performance: the former evaluates out-of-sample discrimination on unseen blocks, while the latter tests whether predicted extremes generalize spatially across corridor elements. The primary limitation remains the absence of event-level crash geolocation and timestamps, which preclude direct external validation of crash matches. Future work can strengthen external validation by leveraging crash datasets with event-level spatial–temporal precision and by implementing explicit spatial cross-validation designs (e.g., leave-one-site-out) when site-level sample sizes become sufficiently large.
@@ 7 Summary and conclusions
This study addresses the challenge of estimating short-term crash occurrence risk (COR) in urban corridors, where conventional surrogate safety approaches often rely on simplified interaction geometry and provide limited support for corridor-scale inference under heterogeneous operations. Existing trajectory-based EVT studies have primarily emphasized vehicle–vehicle (V–V) interactions and rarely incorporate vehicle–infrastructure (V–I) conflicts within a unified corridor-scale framework, thereby limiting their applicability for proactive safety monitoring and infrastructure-relevant risk assessment. To bridge these gaps, we proposed a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator that propagates oriented vehicle footprints and integrates HD-map roadway boundaries, enabling consistent extraction of extreme near-miss events from both V–V and V–I interactions.
@@
Extreme near-misses were modeled using a nonstationary univariate generalized extreme value (UGEV) distribution embedded in a Hierarchical Bayesian Structure Grouped Random Parameters (HBSGRP) framework. The grouped structure supports partial pooling across intersections and directional segments while retaining location-specific heterogeneity, improving stability under sparse extreme samples. High-resolution trajectories from the Argoverse-2 dataset were combined with HD maps to construct interaction-window block maxima of 2D-TTC and to derive covariates representing (i) interaction dynamics (e.g., relative speed, acceleration/deceleration, jerk, and heading/steering differences, with exposure represented by traffic volume) and (ii) roadway context (e.g., lane width, driveway density, median type, turning movements, and lane position). Grouped random parameters were assigned to a small, theoretically motivated subset of high-impact dynamics to capture context-dependent effects without overfitting, and a baseline hierarchical Bayesian fixed-parameter model (HBSFP) was estimated for benchmarking.
Results indicate that the HBSGRP specification consistently improves predictive performance and posterior stability relative to HBSFP, as reflected by lower DIC, WAIC, and LOOIC values. Importantly, covariate effects were interpreted in terms of their implied impact on COR (i.e., on the fitted UGEV distribution and crash-level exceedance probability), rather than through individual GEV parameters in isolation. Under this distribution-level interpretation, closing-dynamics variables (relative speed, deceleration, and distance) emerged as the most influential drivers of crash-prone extreme risk for V–V interactions, particularly on segments, while V–I risk was more strongly tied to spacing/proximity conditions and boundary encroachment. At intersections, V–V near misses were dominated by relative speed and distance, while no dynamic variables were significant for V–I near misses. Roadway and maneuver context (e.g., steering/heading deviations, lane width, driveway density, median type, and turning movements) also contributed meaningfully to COR variation, reinforcing that corridor risk is jointly shaped by short-horizon behavior and geometric/infrastructure constraints.
Model credibility was supported through validation against held-out outcomes using ROC–AUC. Predictive discrimination was strongest for V–V segment models (AUC $\approx$ 0.89), followed by V–V intersection models (AUC $\approx$ 0.82), while V–I models achieved moderate performance (segments: $\approx$ 0.79; intersections: $\approx$ 0.75). These results suggest that V–V crash-prone extremes are more consistently explainable from interaction dynamics at the corridor scale, whereas V–I interactions—particularly at intersections—exhibit greater variability due to complex boundary approach behavior, turning paths, and map-defined infrastructure constraints.
While the proposed framework advances corridor-wide COR estimation, several limitations motivate future research. First, Argoverse-2 motion-forecasting scenarios are intentionally curated toward safety-critical, high-interaction episodes; therefore, the estimated COR should be interpreted as conditional on these traffic states. Second, the 2D-TTC projection assumes simplified short-horizon dynamics, which may underrepresent risk under extreme evasive maneuvers or highly nonlinear control responses. Third, V–V and V–I extremes are modeled as separate univariate processes under the interaction-window block definition; extending to a bivariate or multivariate hierarchical EVT formulation would require a common space–time block that yields paired extremes and sufficient information to identify joint tail dependence. Finally, the current study focuses on crash occurrence risk from extreme near-misses, without explicit severity stratification within the EVT layer.
Several extensions can further strengthen the proposed framework. First, incorporating longer-duration and more representative trajectory streams from continuously operating AV fleets would allow benchmarking the estimated COR against network-wide baseline risk and assessing temporal stability beyond short interaction windows. Second, the current univariate EVT formulation could be extended to multivariate hierarchical EVT models that jointly represent V–V and V–I extremes, provided that a common space–time blocking strategy yields paired extremes with sufficient information to identify tail dependence. Third, integrating complementary physics-informed severity surrogates, such as post-impact delta-V or energy-based indicators, would enable joint modeling of crash occurrence and severity rather than occurrence alone. Fourth, extending the framework to include vulnerable road user (VRU) near-misses would improve relevance for urban safety applications. Finally, systematic comparison with POT–GPD-based formulations under comparable blocking and dependence controls would help clarify trade-offs between BM and POT approaches for short-horizon, high-frequency trajectory data. As trajectory data scale in coverage and duration, hierarchical EVT models such as HBSGRP–UGEV will gain additional statistical power, enabling more nuanced, multimodal safety assessments. Within this evolving data landscape, the proposed framework provides a scalable foundation for proactive corridor-wide monitoring, supporting earlier identification of crash-prone conditions and enabling data-driven traffic management consistent with Vision Zero goals.